OpenStudy (anonymous):

Solve for X by completing the square : x^2 + bx+c=0 thanks !

OpenStudy (nikvist):

\[x^2+bx+c=0\]\[x^2+2\frac{b}{2}x+\left(\frac{b}{2}\right)^2-\left(\frac{b}{2}\right)^2+c=0\]\[\left(x+\frac{b}{2}\right)^2=\left(\frac{b}{2}\right)^2-c\]\[x+\frac{b}{2}=\pm\sqrt{\frac{b^2}{4}-c}=\pm\frac{\sqrt{b^2-4c}}{2}\]\[x=\frac{-b\pm\sqrt{b^2-4c}}{2}\]

OpenStudy (anonymous):

Thank you !